20.0 mm 9.0 x 109 N. m2/C2 Two electrons are passing apart. What is the electric repulsive force that they exert on each other? The value of k is 5.8 x 10 25 N 2.3 N 2.3 x 1010 N 5.8 x 10 27 N
Answer #1

The repulsive force between electrons with distance between these electrons d = 20.0 mm can be calculated as;

\small F=\frac{kq_{e}^{2}}{d^{2}}

\small \Rightarrow F=\frac{9\times 10^{9}\times (1.6\times 10^{-19})^{2}}{(20.0\times 10^{-3})^{2}}

\small \Rightarrow F=5.8\times 10^{-25}\: N